OpenStudy (anonymous):

Find the slope and y-int of each line, then graph. 5) 2x-y=2 6) y=3x+6 ?

OpenStudy (anonymous):

Here I'd start by getting it into slope intercept form. (Get y by itself)

hartnn (hartnn):

for y= mx+c slope = m, y-intercept = c

hartnn (hartnn):

6) is already of the form y=mx+c, so finding m,c is easy. try to isolate y in 5)

OpenStudy (anonymous):

5) y=2x-2 y-int=-2 slope=2